# Learning with less: label-efficient land cover classification at very high spatial resolution using self-supervised deep learning

### v1

@@ 3.4 Model ensembling, large-scale land cover mapping, and validation
Once the best-performing combination of pre-training procedure, semantic segmentation architecture, and training dataset size was determined, we used the models trained using this configuration across all folds to produce a model ensemble for large-scale land cover mapping over the state of Mississippi. All models in the ensemble shared the same architecture, pre-training procedure, and training dataset size, differing only in the specific samples used for training due to the k-fold cross-validation procedure. For example, if the best-performing configuration is found to be a U-Net architecture with BYOL pre-training using 750 training samples, then the final ensemble would consist of 4 U-Net models, each pre-trained using BYOL and fine-tuned using 750 training samples from each fold of the cross-validation procedure, with each model having seen a different subset of the training data. Model ensembling was chosen as it is a popular approach to improving the robustness and accuracy of machine learning models by combining the predictions of multiple distinct models to produce a final prediction. Each model in the ensemble produced a land cover probability vector for each pixel in the input image, and these class probability vectors were averaged across all models in the ensemble to produce a final class probability vector for each pixel, with the final predicted class being the class with the highest probability from the averaged class probabilities [102]. A “sliding window” approach was used to process the imagery in patches with a stride of to ensure that each pixel would be seen multiple times to further maximize the effects of the model ensemble. Additionally, test-time augmentations were employed to further improve the robustness of the predictions: at any given location, each model in the ensemble received an input that was randomly augmented using random horizontal and vertical flips along with random 90∘ rotation [103]. For example, if there are 4 models in the ensemble, each model would receive a different random augmentation of the input patch, resulting in 4 different predictions for that patch. The final output patches of the ensemble were weighted using element-wise multiplication with a 256256 Gaussian kernel to de-emphasize the probabilities of pixels near the edge of the classification maps in order to mitigate presence of edge artifacts in the final merged output. These inference steps were used to classify the entire state of Mississippi using 2023 NAIP imagery at resolution, amounting to over 123 billion pixels. A thematic accuracy assessment of the final land cover product was performed using the point-based assessment dataset (see Section 3.1). Overall accuracy, per-class user accuracy, and per-class producer accuracy metrics were computed to assess the quality of the final product, as shown in Equations 5, 6, and 7, where TP is the number of true positives, FP is the number of false positives, and FN is the number of false negatives.
@@ 4 Results
The evaluation of the self-supervised learning approach is divided into two parts: linear probing performance and fine-tuning results. The linear probing experiments are intended to provide a measure of the quality of the learned image representations produced by the pre-trained encoder in isolation (i.e., without any deep segmentation heads or fine-tuning). As shown in Figure 7, using the embeddings from a ResNet-101 model pre-trained using BYOL as input to a linear classifier yielded substantial improvements in macro F1 score over a classifier that used embeddings from a model pre-trained only on ImageNet. The smallest training dataset size of exhibited the largest improvement in macro F1 score, with the BYOL pre-trained model achieving a macro F1 score of % compared to % for the ImageNet pre-trained model, an improvement of %. The embedding model where BYOL pre-training was initialized with ImageNet weights performed marginally worse than the model that began BYOL pre-training with random weights: the average increase in macro F1 score over the ImageNet pre-trained model was % for the BYOL model with random initialization, compared to % for the BYOL model with ImageNet initialization.
@@
While the linear probing results show vast improvements in representation quality when using BYOL pre-training, fine-tuning these pre-trained encoders using various semantic segmentation architectures resulted in different performance trends. Figure 8 shows that while the BYOL models that were initialized with ImageNet weights prior to BYOL pre-training outperformed ImageNet models with no additional pre-training, the BYOL models that were initialized with random weights yielded more variable performance depending on the specifics of the semantic segmentation architecture. The BYOL encoder with random initialization outperformed the ImageNet pre-trained encoder in terms of mean macro F1 score across the UPerNet and DeepLabV3+ models at , but lagged behind all other architectures at and . Overall, the ImageNet pre-trained models had a mean macro F1 score of %, while the BYOL model with random initialization had a mean macro F1 score of %. However, the BYOL model initialized with ImageNet weights prior to pre-training consistently outperformed the ImageNet pre-trained model across all architectures and training dataset sizes, achieving a mean macro F1 score of %.
@@
Increasing the training dataset size generally led to improvements in macro F1 score across all architectures and pre-training strategies, though the magnitude of these improvements diminished as the dataset size increased. For example, the average increase in macro F1 score when increasing the training dataset size from to was %, while the average increase from to was only %. The impact of increasing training dataset size was most pronounced for the ImageNet pre-trained models, which saw a mean increase of % in macro F1 score when increasing the training dataset size from to , compared to only % for the BYOL (ImageNet initialization) models and % for the BYOL (random initialization) models. In terms of semantic segmentation architecture, the U-Net models had the highest mean macro F1 score across all training dataset sizes and pre-training strategies (71.20%), followed by Attention U-Net (70.85%), UPerNet (69.01%), DeepLabV3+ (67.96%), PAN (67.32%), and FCN (65.22%). No correlation between model size (in terms of number of parameters) and macro F1 score was observed (), indicating that there was no clear relationship between model complexity and performance in this context.
@@
The best-performing pre-training, semantic segmentation architecture, and training dataset size combination based on mean macro F1 score across all k-folds was the BYOL model initialized with ImageNet weights used as the backbone in a U-Net architecture trained with samples, which achieved a mean macro F1 score of % and a mean overall accuracy of %, a % and % improvement over the U-Net model using the same training dataset size but with an ImageNet pre-trained backbone, respectively. To illustrate the qualitative benefits of the BYOL pre-training approach, we highlight a particularly strong individual model run. Run #2 of the U-Net model with the BYOL backbone (ImageNet initialization) fine-tuned using samples achieved a macro F1 score of %. A qualitative assessment of this model’s inference results is shown in Figure 9, demonstrating that not only did the model outperform its ImageNet-only counterpart quantitatively (an increase of % macro F1 score for the same data split), but it also produced land cover maps with sharper object boundaries and fewer classification artifacts. Samples (a), (b), and (c) in Figure 9 show overall improvements in classification accuracy with additional pre-training across multiple land cover classes, particularly in homogeneous regions such as water bodies and fields (though some misclassifications still occur, such as in (c)). The importance of object localization is highlighted in samples (d), (e), and (f), where the ImageNet-only model tends to yield acceptable overall classification accuracy, but fails to accurately delineate small features such as roads and buildings, which are better captured by the model with additional BYOL pre-training. While these examples are illustrative and do not change the overall ranking reported above, they highlight the qualitative improvements of the BYOL pre-training approach, which are particularly important given the spatial resolution of the ground truth labels.
@@
Despite this specific model’s large individual performance, it was an outlier in terms of performance across all k-folds, as the mean macro F1 score and overall accuracy for this model were % and %, respectively, falling short of the best-performing configuration in terms of mean macro F1 score. As such, the U-Net model with a BYOL backbone initialized with ImageNet weights and fine-tuned using samples was considered the best-performing model overall and was used to generate the final land cover map for the entire state of Mississippi. The 4 models trained using this best-performing configuration as part of the k-fold cross-validation procedure were subsequently used to generate a final land cover map for the entire state of Mississippi using the model ensembling and test-time augmentation procedure outlined in Section 3.4. A visualization of the final comprehensive resolution land cover map for Mississippi is shown in Figure 10. We note the presence of a large vertical stripe of barren land in the Mississippi Delta region that stands in contrast to the surrounding cropland; this is an artifact of a data acquisition anomaly in the 2023 NAIP imagery, where a portion of the imagery was collected during a period when fields were fallow and thus appeared barren. This highlights a unique challenge of mapping agricultural activity using imagery acquired across multiple dates, as phenological changes in cropland can cause inconsistent land cover classifications. Still, despite this anomaly, our final land cover product offers a complete and precise spatial inventory of natural resources and human developments across the entire state. For example, water bodies appear to be delineated with high precision, revealing the presence of small streams and ponds, forested areas are well captured with fragmentation appropriate to the landscape context (e.g., trees interspersed within urban developments), and developed areas with small buildings and road networks are distinct and mapped with adequate spatial detail.
@@
The confusion matrix from final accuracy assessment is shown in Table 3, demonstrating that our model yields high accuracy when classifying natural resources, while struggling with land cover types that are more anthropogenically influenced. For example, the barren land class has a moderately high user’s accuracy of %, but a low producer’s accuracy of %, and the model has a tendency to misclassify barren land pixels as impervious surfaces, herbaceous, and cultivated crops. The cultivated crops class suffers from the opposite problem, with a high producer’s accuracy of % but a low user’s accuracy of %, with barren land and herbaceous classes being frequently misclassified as cultivated crops. On the other hand, the open water and tree canopy/woody vegetation classes are classified with high accuracy, with F1 scores of % and %, respectively, highlighting the model’s strength in identifying these natural features. Overall, the final product’s macro F1 score and overall accuracy were % and %, respectively, while the mean class-wise producer’s accuracy and user’s accuracy were % and %, respectively. We do note that the large amount of forested land in Mississippi makes a balanced assessment challenging, as forested land accounted for over % of all sampled assessment points, making the overall accuracy metric less informative due to the class imbalance: we thus emphasize the macro F1 score and individual class-wise metrics as a more holistic measure of model performance in this context.
| Open | |||||||||
| Water | Impervious | ||||||||
| Structures | Impervious | ||||||||
| Surfaces | Barren | ||||||||
| Land | Tree Canopy/ | ||||||||
| Woody Vegetation | Low Vegetation/ | ||||||||
| Herbaceous | Cultivated | ||||||||
| Crops | Unclassified | Producer’s | |||||||
| Accuracy | |||||||||
| Open Water | 511 | 0 | 10 | 11 | 18 | 0 | 21 | 89.34% | |
| Impervious | |||||||||
| Structures | 0 | 15 | 3 | 3 | 5 | 0 | 0 | 75.47% | |
| Impervious | |||||||||
| Surfaces | 1 | 183 | 17 | 1 | 22 | 0 | 1 | 79.57% | |
| Barren Land | 6 | 68 | 944 | 22 | 451 | 144 | 5 | 57.42% | |
| Tree Canopy/ | |||||||||
| Woody Vegetation | 1 | 2 | 6 | 14712 | 366 | 6 | 22 | 97.32% | |
| Low Vegetation/ | |||||||||
| Herbaceous | 7 | 20 | 138 | 376 | 3699 | 492 | 17 | 77.86% | |
| Cultivated Crops | 0 | 0 | 17 | 2 | 38 | 1243 | 1 | 95.54% | |
| Unclassified | 2 | 6 | 7 | 753 | 86 | 1 | 413 | 32.29% | |
| User’s Accuracy | 96.78% | 62.24% | 82.66% | 92.64% | 78.95% | 65.91% | 86.04% | Overall Accuracy | |
| 87.14% | |||||||||
@@
The final land cover product’s macro F1 score represents an increase of % over the mean performance of the base k-fold models, which can be attributed to the model ensembling and test-time augmentation procedures outlined in Section 3.4. Table 4 provides a class-wise comparison of the F1 scores between the performance of the models in the k-fold cross-validation and their performance when ensembled. Notably, the impervious surfaces and barren land classes saw dramatic improvements in F1 score after ensembling, with increases of % and %, respectively. Other classes saw more modest improvements, while the unclassified class saw a decrease of % in F1 score after ensembling. The use of the model ensemble approximately quadrupled the time needed for inference compared to a single model, highlighting the trade-off between accuracy and computational efficiency when using model ensembling and test-time augmentation. For our purposes, we deemed the increase in accuracy worth the additional computational cost, which was not prohibitive given the available computational resources, and elected to use the ensembled predictions for the final land cover product.
| Open | ||||||||
| Water | Impervious | |||||||
| Structures | Impervious | |||||||
| Surfaces | Barren | |||||||
| Land | Tree Canopy/ | |||||||
| Woody Vegetation | Low Vegetation/ | |||||||
| Herbaceous | Cultivated | |||||||
| Crops | Unclassified | |||||||
| Mean Ensemble F1 | 92.24 | 63.39 | 63.96 | 94.84 | 77.01 | 76.55 | 49.04 | |
| Final F1 | 92.91 | 69.85 | 67.77 | 94.93 | 78.40 | 78.00 | 46.96 | |
| Improvement | +0.67 | +6.46 | +3.81 | +0.09 | +1.40 | +1.45 | -2.09 |

### v2

@@ 3.4 Model ensembling, large-scale land cover mapping, and validation
Once the best-performing combination of pre-training procedure, semantic segmentation architecture, and training dataset size was determined, we used the models trained using this configuration across all folds to produce a model ensemble for large-scale land cover mapping over the state of Mississippi. All models in the ensemble shared the same architecture, pre-training procedure, and training dataset size, differing only in the specific samples used for training due to the k-fold cross-validation procedure. For example, if the best-performing configuration is found to be a U-Net architecture with BYOL pre-training using 750 training samples, then the final ensemble would consist of 4 U-Net models, each pre-trained using BYOL and fine-tuned using 750 training samples from each fold of the cross-validation procedure, with each model having seen a different subset of the training data. Model ensembling was chosen as it is a popular approach to improving the robustness and accuracy of machine learning models by combining the predictions of multiple distinct models to produce a final prediction. Each model in the ensemble produced a land cover probability vector for each pixel in the input image, and these class probability vectors were averaged across all models in the ensemble to produce a final class probability vector for each pixel, with the final predicted class being the class with the highest probability from the averaged class probabilities [84]. A “sliding window” approach was used to process the imagery in patches with a stride of to ensure that each pixel would be seen multiple times to further maximize the effects of the model ensemble. Additionally, test-time augmentations were employed to further improve the robustness of the predictions: at any given location, each model in the ensemble received an input that was randomly augmented using random horizontal and vertical flips along with random 90∘ rotation [91]. For example, if there are 4 models in the ensemble, each model would receive a different random augmentation of the input patch, resulting in 4 different predictions for that patch. The final output patches of the ensemble were weighted using element-wise multiplication with a 256256 Gaussian kernel to de-emphasize the probabilities of pixels near the edge of the classification maps in order to mitigate presence of edge artifacts in the final merged output. These inference steps were used to classify the entire state of Mississippi using 2023 NAIP imagery at resolution, amounting to over 123 billion pixels. A thematic accuracy assessment of the final land cover product was performed using the point-based assessment dataset (see Section 3.1). Overall accuracy, user’s accuracy, producer’s accuracy, F1 Score, Intersection-over-Union (IoU), and Cohen’s Kappa () metrics were computed to assess the quality of the final product, as shown in Equations 6–10, where TP is the number of true positives, FP is the number of false positives, and FN is the number of false negatives [23, 89].
3.5 Cross-year generalization and validation
To examine the capacity of the model to generalize to other years for the purpose of creating multi-temporal land cover/land cover change products, we applied the final model ensemble to 2016 NAIP imagery over Mississippi to create a secondary land cover product using the same protocol outlined in Section 3.4. Beyond the creation of a multi-temporal land cover data product, this also enabled an evaluation of the model’s capability of handling changes in land cover distribution. Additionally, NAIP acquisition characteristics vary between years, in terms of both sensor specifications and acquisition period, making it a suitable tool for evaluating the model’s response to domain shifts. As NAIP is a basic aerial imagery product (and does not provide calibrated reflectance values), we adjusted the 2016 imagery to match the spectral characteristics of the 2023 imagery using histogram normalization. The 2016 product was validated using the same locations as the point-based assessment dataset for 2023 (25,000 samples), but with new human-annotated ground truth labels that correspond to 2016 land cover.
@@ 4 Results
The evaluation of the self-supervised learning approach is divided into two parts: linear probing performance and fine-tuning results. The linear probing experiments are intended to provide a measure of the quality of the learned image representations produced by the pre-trained encoder in isolation (i.e., without any deep segmentation heads or fine-tuning). As shown in Figure 7, using the embeddings from a ResNet-101 model pre-trained using BYOL only as input to a linear classifier yielded an average 15.44% improvement in macro F1 score over a classifier that used embeddings from a model pre-trained only on ImageNet, which outperformed the models pre-trained only with MoCoV2 in terms of macro F1 score by 23.48% on average. This indicates that the BYOL successfully enabled a model to learn to extract features that are linearly separable with regard to land cover classification without the need for ImageNet initialization, while MoCoV2 failed to do so.
Initializing the ResNet-101 model with ImageNet weights prior to MoCoV2 pre-training yielded moderately better performance, resulting in a 9.59% average increase in macro F1 score compared to simply starting from ImageNet alone, but the BYOL models with ImageNet pre-training outperformed these MoCoV2 models by 5.27%. The smallest training dataset size of exhibited the largest improvement in macro F1 score over ImageNet initialization from BYOL pre-training, with the BYOL pre-trained models achieving a mean macro F1 score of 51.78% compared to 35.01% for the ImageNet pre-trained model, an improvement of 16.77%. Notably, the embedding model where BYOL pre-training was initialized with ImageNet weights performed marginally worse than the model where BYOL pre-training was performed with random initialization: the average increase in macro F1 score over the ImageNet pre-trained model was 15.44% for the BYOL model with random initialization, compared to 14.87% for the BYOL model with ImageNet initialization, a 0.58% difference. Overall, the BYOL models showed clear superiority over the MoCoV2 models and the ImageNet model (regardless of initialization strategy), while the MoCoV2 initialized with ImageNet weights showed moderate feature extraction capability compared to an ImageNet model.
@@
While the linear probing results show vast improvements in representation quality when using BYOL pre-training, fine-tuning these pre-trained encoders using various semantic segmentation architectures resulted in different performance trends. Figure 8 shows that the ImageNet encoder outperformed both the BYOL and MoCoV2 pre-trained encoders with random initialization across all combinations of segmentation architectures, with ImageNet models averaging 70.10% in terms of macro F1 score, a 1.53% improvement over BYOL and a 5.70% over MoCoV2. These results show that, while not achieving the same performance as ImageNet fully supervised pre-training, utilizing self-supervised pre-training with BYOL on domain specific data with randomly initialized weights can be nearly as effective. Additionally, BYOL with random initialization consistently outperformed MoCoV2 with random initialization across all experimental axes, with an average improvement in macro F1 score of 4.17%. MoCoV2 with random initialization also exhibited large sensitivity to shifting training data during cross-validation with an overall standard deviation of 3.39%, more than double that of BYOL with random initialization (1.52%) or ImageNet (1.48%).
Initializing models with ImageNet weights prior to pre-training proved to yield the highest performance: our BYOL and MoCoV2 models pre-trained using this approach achieved average macro F1 scores of 71.63% and 70.62%, respectively, increases of 1.53% and 0.52% over the ImageNet baseline. The BYOL models initialized with ImageNet prior to pre-training outperformed all ImageNet models across all test cases, while MoCoV2 outperformed the ImageNet encoder across all cases with the exception of the UPerNet networks. BYOL with ImageNet pre-training also outperformed MoCoV2 pre-training in all but one case, where the FCN model trained with 750 training samples outperformed its BYOL counterpart by 0.16%. BYOL with ImageNet initialization also demonstrated the least variation across all encoder initialization frameworks, with a standard deviation of 1.22% compared to ImageNet’s 1.48% and ImageNet followed by MoCoV2 pre-training’s 1.66%.
@@
Increasing the training dataset size generally led to improvements in macro F1 score across all architectures and pre-training strategies, though the magnitude of these improvements diminished as the dataset size increased. For example, the average increase in macro F1 score when increasing the training dataset size from to was 2.62%, while the average increase from to was only 0.69%. The impact of increasing training dataset size was most pronounced for the MoCoV2 pre-trained models with random initialization, which saw a mean increase of 6.45% in macro F1 score when increasing the training dataset size from to , followed by 2.98% for the ImageNet models, 2.68% for the MoCoV2 models with ImageNet initialization, 2.45% for the BYOL models with random initialization, and 2.02% for the BYOL models with ImageNet initialization. In terms of semantic segmentation architecture, the U-Net models had the highest mean macro F1 score across all training dataset sizes and pre-training strategies (71.48%), followed by Attention U-Net (71.06%), UPerNet (69.16%), DeepLabV3+ (68.80%), PAN (68.07%), and FCN (65.79%). Weak positive correlation between model size (in terms of number of parameters) and macro F1 score was observed (). A comprehensive table of performance metrics across all models tested is provided in Table S1 in the supplementary materials.
@@
The best-performing pre-training, semantic segmentation architecture, and training dataset size combination based on mean macro F1 score across all k-folds was the BYOL model initialized with ImageNet weights used as the backbone in a U-Net architecture trained with samples, which achieved a mean macro F1 score of 74.04% and a mean overall accuracy of 86.43%, a 1.47% and 0.32% improvement over the U-Net model using the same training dataset size but with an ImageNet pre-trained backbone, respectively. To illustrate the qualitative benefits of the BYOL pre-training approach, we highlight a particularly strong individual model run. Run #2 of the U-Net model with the BYOL backbone (ImageNet initialization) fine-tuned using samples achieved a macro F1 score of 75.84%. A qualitative assessment of this model’s inference results is shown in Figure 9, demonstrating that not only did the model outperform its ImageNet-only counterpart quantitatively (an increase of 2.06% macro F1 score for the same data split), but it also produced land cover maps with sharper object boundaries and fewer classification artifacts. Samples (a), (b), and (c) in Figure 9 show overall improvements in classification accuracy with additional pre-training across multiple land cover classes, particularly in homogeneous regions such as water bodies and fields (though some misclassifications still occur, such as in (c)). The importance of object localization is highlighted in samples (d), (e), and (f), where the ImageNet-only model tends to yield acceptable overall classification accuracy, but fails to accurately delineate small features such as roads and buildings, which are better captured by the model with additional BYOL pre-training. While these examples are illustrative and do not change the overall ranking reported above, they highlight the qualitative improvements of the BYOL pre-training approach, which are particularly important given the spatial resolution of the ground truth labels.
@@
Despite this specific model’s strong individual performance, it was an outlier in terms of performance across all k-folds, as the mean macro F1 score and overall accuracy for this model were 73.70% and 86.55%, respectively, falling short of the best-performing configuration in terms of mean macro F1 score. As such, the U-Net model with a BYOL backbone initialized with ImageNet weights and fine-tuned using samples was considered the best-performing model overall and was used to generate the final land cover map for the entire state of Mississippi. The 4 models trained using this best-performing configuration as part of the k-fold cross-validation procedure were subsequently used to generate a final land cover map for the entire state of Mississippi using the model ensembling and test-time augmentation procedure outlined in Section 3.4. A visualization of the final comprehensive resolution land cover map for Mississippi is shown in Figure 10. We note the presence of a large vertical stripe of barren land in the Mississippi Delta region that stands in contrast to the surrounding cropland; this is an artifact of a data acquisition anomaly in the 2023 NAIP imagery, where a portion of the imagery was collected during a period when fields were fallow and thus appeared barren. This highlights a unique challenge of mapping agricultural activity using imagery acquired across multiple dates, as phenological changes in cropland can cause inconsistent land cover classifications. Still, despite this anomaly, our final land cover product offers a complete and precise spatial inventory of natural resources and human developments across the entire state. For example, water bodies appear to be delineated with high precision, revealing the presence of small streams and ponds, forested areas are well captured with fragmentation appropriate to the landscape context (e.g., trees interspersed within urban developments), and developed areas with small buildings and road networks are distinct and mapped with adequate spatial detail.
@@
The confusion matrix from final accuracy assessment is shown in Table 3, while Table 4 gives a comprehensive breakdown of the model’s per-class and overall accuracy metrics, demonstrating that our model yielded high accuracy when classifying natural resources, but struggled with land cover types that are more anthropogenically influenced. For example, the barren land class had a moderately high user’s accuracy of 82.66%, but a low producer’s accuracy of 57.42%, and the model had a tendency to misclassify barren land pixels as impervious surfaces, herbaceous, and cultivated crops. The cultivated crops class suffered from the opposite problem, with a high producer’s accuracy of 95.54% but a low user’s accuracy of 65.91%, with barren land and herbaceous classes being frequently misclassified as cultivated crops. On the other hand, the open water and tree canopy/woody vegetation classes were classified with high accuracy, with F1 scores of 92.91% and 94.93%, respectively, highlighting the model’s strength in identifying these natural features. Overall, the final product’s macro F1 score and overall accuracy were 75.58% and 87.14%, respectively, while the mean class-wise user’s and producer’s accuracy were 80.18% and 75.60%, respectively. We do note that the large amount of forested land in Mississippi makes a balanced assessment challenging, as forested land accounted for over 60% of all sampled assessment points, making the overall accuracy metric less informative due to the class imbalance: we thus emphasize the macro F1 score and individual class-wise metrics as a more holistic measure of model performance in this context. A thorough quantitative and qualitative assessment of errors is given in Section S2 of the supplementary materials.
| Predicted Class | |||||||||
| True Class | Open Water | Impervious Structures | Impervious Surfaces | Barren Land | Tree Canopy/Woody Vegetation | Low Vegetation/ Herbaceous | Cultivated Crops | Unclassified | Total |
| Open Water | 511 | 1 | 0 | 10 | 11 | 18 | 0 | 21 | 572 |
| Impervious Structures | 0 | 80 | 15 | 3 | 3 | 5 | 0 | 0 | 106 |
| Impervious Surfaces | 1 | 5 | 183 | 17 | 1 | 22 | 0 | 1 | 230 |
| Barren Land | 6 | 4 | 68 | 944 | 22 | 451 | 144 | 5 | 1,644 |
| Tree Canopy | 1 | 2 | 2 | 6 | 14,712 | 366 | 6 | 22 | 15,117 |
| Low Vegetation | 7 | 2 | 20 | 138 | 376 | 3,699 | 492 | 17 | 4,751 |
| Cultivated Crops | 0 | 0 | 0 | 17 | 2 | 38 | 1,243 | 1 | 1,301 |
| Unclassified | 2 | 11 | 6 | 7 | 753 | 86 | 1 | 413 | 1,279 |
| Total | 528 | 105 | 294 | 1,142 | 15,880 | 4,685 | 1,886 | 480 | 25,000 |
| Class Name | User’s Accuracy (%) | Producer’s Accuracy (%) | F1 Score (%) | IoU (%) | Cohen’s |
| Open Water | 96.78 | 89.34 | 92.91 | 86.76 | 0.9275 |
| Impervious Structures | 76.19 | 75.47 | 75.83 | 61.07 | 0.7573 |
| Impervious Surfaces | 62.24 | 79.57 | 69.85 | 53.67 | 0.6953 |
| Barren Land | 82.66 | 57.42 | 67.77 | 51.25 | 0.6593 |
| Tree Canopy / Woody Veg. | 92.64 | 97.32 | 94.93 | 90.34 | 0.8666 |
| Low Veg. / Herbaceous | 78.95 | 77.86 | 78.40 | 64.48 | 0.7338 |
| Cultivated Crops | 65.91 | 95.54 | 78.00 | 63.94 | 0.7656 |
| Unclassified | 86.04 | 32.29 | 46.96 | 30.68 | 0.4543 |
| Macro Average | 80.18 | 75.60 | 75.58 | 62.77 | 0.7325 |
| Overall | Accuracy: 87.14 | 77.21 | 0.7751 |
@@
The final land cover product’s macro F1 score represents an increase of 1.54% over the mean performance of the base k-fold models, which can be attributed to the model ensembling and test-time augmentation procedures outlined in Section 3.4. Table 5 provides a class-wise comparison of the F1 scores between the best overall performing model, the mean performance of the models in the ensemble, the best performing model in the ensemble, and the model ensemble itself. Notably, the impervious surfaces and barren land classes saw dramatic improvements in F1 score compared to the mean performance of the models in the ensemble, with increases of 6.46% and 3.81%, respectively. Other classes saw more modest improvements, while the unclassified class showed a decrease of 2.09% in F1 score after ensembling. The best performing model in the ensemble (from cross-validation run #2) outperformed the ensemble itself when classifying impervious structures (+1.09% F1 score), but was more than offset by the ensemble’s capacity for classifying impervious surfaces (+5.51% F1 score), leading to an improvement of 0.58% macro F1 score and 0.53% overall accuracy.
Overall | ||||||||||||||||||
92.74 | 75.98 | 62.74 | 68.10 | 95.12 | 78.11 | 81.23 | 52.73 | 75.84 | 87.42 | |||||||||||||||||
92.24 | 75.28 | 63.39 | 63.96 | 94.84 | 77.01 | 76.55 | 49.04 | 74.04 | 86.43 | |||||||||||||||||
92.66 | 76.92 | 64.34 | 67.07 | 94.88 | 76.92 | 76.34 | 50.84 | 75.00 | 86.61 | |||||||||||||||||
92.91 | 75.83 | 69.85 | 67.77 | 94.93 | 78.40 | 78.00 | 46.96 | 75.58 | 87.14 |
When compared to the best performing model overall (U-Net, ImageNet + BYOL backbone, , cross-validation run #2), we observe that the ensemble performed similarly on aggregate, with a marginal 0.26% decrease in macro F1 score and a 0.28% decrease in overall accuracy, but the ensemble yielded a 7.11% increase in F1 score in the classification of impervious surfaces, a critical minority class for VHSR land cover data. The best individual model’s primary advantage over the ensemble was the classification of “unclassified” regions, with an improvement of 5.77% in terms of F1 score. However, as the unclassified class provides little utility for downstream analysis, we selected the ensemble model for its improved accuracy in the impervious surface class for robust mapping of urban regions. The use of the model ensemble approximately quadrupled the time needed for inference compared to a single model, highlighting the trade-off between accuracy and computational efficiency when using model ensembling and test-time augmentation. For our purposes, we deemed the increase in accuracy worth the additional computational cost, which was not prohibitive given the available computational resources, and elected to use the ensembled predictions for the final land cover product.
As shown in Table 6, the model was able to classify 2016 imagery well despite not being trained on this data. Only relatively small decreases of 2.36% and 1.93% were observed in overall accuracy and macro F1 score. However, there was considerable variation in class-wise performance between the two years: for example, the 2016 product yielded a 5.73% improvement in F1 score over the 2023 product when classifying barren land classes, but showed a 9.05% decrease in F1 score when classifying cultivated crops. One potential cause for drastic shift in classification accuracy is the difference in acquisition schedule for the 2016 imagery compared to the 2023 imagery. Most of the imagery for 2016 was acquired in early June (either during or shortly after the planting period for soybeans [114]) whereas a majority of the 2023 imagery was acquired in mid-to-late August. Figure 12 shows that the model predicted 28.92% less cropland in 2016 compared to 2023, while the predicted barren land area increased by 41.24%. Examining the spatial distribution of the 2016 land cover data shows that most of the larger swaths of barren land lie in the Mississippi Delta region of the state interleaved with other cultivated crops. This highlights the unique challenge of mapping agricultural activity from mono-temporal aerial imagery, particularly in Mississippi where a wide range of crops with varying phenological characteristics are grown, making it difficult to capture agricultural activity comprehensively. Still, the model performed similarly on the 2016 data as the 2023 data across all other classes, with the exception of the unclassified class. We provide a confusion matrix in Table S2 of the supplementary materials, along with class-wise accuracy metrics in Table S3.
Unclassified |
Overall | |||||||||||||||||
| 2023 F1 Score (%) | 92.91 | 75.83 | 69.85 | 67.77 | 94.93 | 78.40 | 78.00 | 46.96 | 75.58 | 87.14 | ||||||||||||||||
| 2016 F1 Score (%) | 90.26 | 76.53 | 71.82 | 73.49 | 93.26 | 77.18 | 68.96 | 37.70 | 73.65 | 84.78 |